# A new HOC scheme for computation of flow and heat transfer in nonuniform polar grids

### v2

@@ 4.4 Problem 4: Forced convection over a stationary heated circular cylinder
Finally, we carry out numerical simulation for the classical problem of heat transfer in the fluid flow over the stationary heated circular cylinder. The stationary heated circular cylinder is presumed to be of unit radius ($R_{0}=1$) and immersed in a fluid of infinite domain maintained at unit nondimensional temperature [28, 29, 30, 31, 8, 32, 33, 34, 35, 17, 36, 37, 38]. The cylinder is kept in a cross-flow having uniform velocity $u=U_{\infty}=1$. The 2D flow configuration of the problem is shown in fig. 14a. The cylinder is placed at the center of the circular domain. Following [39], we have set the far field boundary at a distance $R_{\infty}=45R_{0}$. On the solid surface $r=R_{0}$, the boundary conditions for velocity components are those of no-slip conditions, i.e. $u=v=0\implies\psi=0$. In the far stream $r=R_{\infty}$ in front of the cylinder the potential flow is prescribed a unit value $u=U_{\infty}$=1. For this problem, where the simulation results in periodic vortex shedding, at the downstream of the flow, we have imposed the convective boundary conditions $\phi_{t}+U_{\infty}\phi_{r}$=0 (where $\phi$ represents $\psi$, $u$ or $v$) to capture the shedding process efficiently. The continuous shedding of vortices, when they leave the computational domain in the direction of the flow, can be best facilitated by the convective boundary conditions [40, 41, 42, 43]. In addition, at the far field vorticity value decays and becomes $\omega=0$. However, it possesses nonzero value at the inner boundary, which may be derived by making use of the fact that $\psi=0$, $\psi_{r}=0$ on $r=R_{0}$ in equation 33b. The vorticity gradients at all the boundaries are computed using one-sided approximations.
@@
d
e
f
Figure 16: Problem 4: Steady state streamlines (left), vortifity contours (middle) and isotherms for (a) $Re$=10, (b) $Re$=20 and (c) $Re$=40.
@@
Present results of vorticity distribution on the cylinder surface for $Re$=10, 20 and 40 and local Nusselt number distribution for $Re$=10, 20 and 45 are portrayed in fig. 17 and 17 respectively; these values are compared with the previous numerical results from Dennis and Chang [44] and Bharti et al. [28]. It can be seen that the two sets of data show excellent agreement with each other. The largest value of the local Nusselt number is obtained at the front stagnation point of the cylinder; the value of the Nusselt number then reduces gradually towards the rear stagnation point and attains its lowest value thereat (see fig. 17). In table 8, we have compiled our computed values of $L/D$, where $L$ is the length of the recirculation bubbled formed behind the cylinder, the drag coefficient $C_{D}$ and the average Nusselt number $\mkern 1.5mu\overline{\mkern-1.5muNu\mkern-1.5mu}\mkern 1.5mu$ for $Re$=10, 20 and 40. As can be seen from the figures and table, the length of the recirculation region and the average Nusselt number depend directly on the value of the Reynolds number, while the drag coefficient varies inversely. A quantitative comparison is also carried out in table 8. Computational results are compared with reference data [28, 29, 30, 31, 8, 32, 33, 34, 35, 17, 36, 37, 38]. Once again, the efficiency of the present scheme is evident from the proximity of our results with the well-established results available in the literature.
@@
Next, we are interested in studying the periodic flow for $Re$=100, 140, and 200. The transition of the vorticity and isotherm contours from steady to periodic state for $Re$=100 is depicted in fig. 18. This may be considered as the representation of the flow for the other $Re$ values considered, although each requires a different time to attain periodicity. As can be seen in fig. 18, no vortex structure formed in the flow field, and the heated cylinder generates a thermal boundary layer near its surface at the beginning. As time marches, two opposite symmetrical vortices are formed simultaneously which grow in width and length with time, and remain symmetrically stable for a certain period as shown in figs. 18 and 18. During this period, the thickness of the thermal boundary layer increases; the temperature distribution also remains symmetric about the line $\theta=0$. After a certain point of time, a fluctuation develops in the flow which destroys the symmetry in the vortices and temperature distribution and they start to oscillate downstream which can be noticed in fig. 18. As the flow develops further, the vortices at the rear of the cylinder grow in size and ultimately are shed from the cylinder towards downstream while heat is carried away. It is heartening to see that the present scheme could capture the characteristic feature of the periodic flow, the so-called von Kármán vortex street very efficiently. With the vortex shedding taking place, isolated hot fluid clusters are observed in the flow (see fig. 18). Subsequently, figs. 18 and 18 indicate the synchronized variation of flow and temperature fields, where vortex shedding phenomena play a determinant role in the heat transfer at the downstream.
@@
The time history of the drag coefficient $(C_{D})$, lift coefficient $(C_{l})$ and average Nusselt number $(\mkern 1.5mu\overline{\mkern-1.5muNu\mkern-1.5mu}\mkern 1.5mu)$ displayed in figs. 19, 19 and 19 pertinently implies that the newly developed scheme has accurately captured the periodic state for all the Reynolds numbers under consideration. The periodicity is further justified by the single dominating peak for $C_{D}$ and $C_{L}$ in the spectral density analysis (see fig. 19). The distribution of local Nusselt number $(Nu^{0})$ over the surface of the cylinder for different $Re$ values are represented in fig. 20. In fig. 20 we carry out a qualitative comparison between the surface distribution of $Nu^{0}$ for $Re$=200 computed in the present study with those of [45, 46]; which reveal close agreement of our results with those taken from the literature. We then compare our numerical values of the flow parameters $C_{d}$, $C_{l}$, $St$ and $\mkern 1.5mu\overline{\mkern-1.5muNu\mkern-1.5mu}\mkern 1.5mu$ with existing studies in table 9. The streakline from our simulation depicting the well-known von Kármán vortex street for $Re$=140 is presented in LABEL:{4fig:P5_streak140} along with the streakline reported in the experimental work of Taneda [47]. From the comparison we note that vortex creation and dissipation as captured in the experimental work is effectively computed by the current formulation. Striking similarity between our numerical results and the existing numerical and experimental results for all the cases indeed verify and validate the scheme.
@@ 5 Conclusion
In this work, we extend the philosophy of recent work on 2D transient convection and diffusion equation on Cartesian coordinates [18] to polar coordinates to tackle non-rectangular geometries. The transformation-free higher-order compact finite difference scheme proposed here can be applied directly on polar grids, unlike most of the schemes in the literature on polar grids that use transformation between physical and computational domains. We have also combined the virtue of nonuniformity with the present scheme. This levitates the efficiency and robustness of the scheme as it acquires the advantage of accumulating and scattering grid points as necessary. The scheme is theoretically third-order accurate in space and second-order accurate temporally. It is worthwhile to mention that for linear problem with analytical solution the scheme exhibits a spatial convergence of order four, which is higher than the theoretical value. The robustness of the scheme is examined by applying it to as many as three benchmark problems of fluid flow and heat flow, viz. driven polar cavity, natural convection in a horizontal concentric annulus, and forced convection around a heated stationary cylinder. In order to resolve the Neumann type boundary conditions, we have also introduced one-sided approximations for first order derivatives. For the problem of forced convection around a heated stationary cylinder, both steady and periodic solutions are being accurately captured by the present scheme. Additionally, important aspects such as the von Kármán vortex phenomenon and the effect of vortex shedding on heat transfer are also studied comprehensively for this fluid body interaction problem. The accuracy of the computed solutions is estimated from the results obtained in all the test problems, which are in excellent agreement with the existing results both qualitatively and quantitatively. We have reported the perceived order of convergence for all the variables in flow problems that are not supplemented with analytical solutions. It is heartening to see that our solutions could attain the theoretical order of convergence in both space and time in all cases.

### v3

@@ 4.4 Problem 4: Forced convection over a stationary heated circular cylinder
Finally, we carry out numerical simulation for the classical problem of heat transfer in the fluid flow over the stationary heated circular cylinder. The stationary heated circular cylinder is presumed to be of unit radius ($R_{0}=1$) and immersed in a fluid of infinite domain maintained at unit nondimensional temperature [30, 31, 33, 37, 34, 36, 28, 35, 8, 38, 17, 29, 32]. The cylinder is kept in a cross-flow having uniform velocity $u=U_{\infty}=1$. The 2D flow configuration of the problem is shown in fig. 13a. The cylinder is placed at the center of the circular domain. Following [39], we have set the far field boundary at a distance $R_{\infty}=45R_{0}$. At the solid surface $r=R_{0}$, the boundary conditions for velocity components are those of no-slip type, i.e. $u=v=0\implies\psi=0$. In the far stream $r=R_{\infty}$, in front of the cylinder the potential flow is prescribed a unit value $u=U_{\infty}$=1. For this problem, where the simulation results in periodic vortex shedding, at the downstream of the flow, we have imposed the convective boundary conditions $\phi_{t}+U_{\infty}\phi_{r}$=0 (where $\phi$ represents $\psi$, $u$ or $v$) to capture the shedding process efficiently. The continuous shedding of vortices, when they leave the computational domain in the direction of the flow, can be best facilitated by the convective boundary conditions [40, 41, 42, 43]. In addition, at the far field vorticity value decays and becomes $\omega=0$. However, it possesses nonzero value at the inner boundary, which may be derived by making use of the fact that $\psi=0$, $\psi_{r}=0$ on $r=R_{0}$ in equation 36b. The vorticity gradients at all the boundaries are computed using one-sided approximations.
@@
Figure 15: Problem 4: Steady state vortifity contours (left) and isotherms (right) for (a) $Re=10$, (b) $Re=20$ and (c) $Re=40$.
@@
Present results of vorticity distribution on the cylinder surface for $Re=10$, 20 and 40 and local Nusselt number distribution for $Re=10$, 20 and 45 are portrayed in fig. 16 and 16 respectively; these values are compared with the previous numerical results from Dennis and Chang [44] and Bharti et al. [28]. It can be seen that the two sets of data show excellent agreement with each other. The largest value of the local Nusselt number is obtained at the front stagnation point of the cylinder; the value of the Nusselt number then reduces gradually towards the rear stagnation point and attains its lowest value thereat (see fig. 16). In table 8, we have compiled our computed values of $L/D$, where $L$ is the length of the recirculation bubbled formed behind the cylinder, the drag coefficient $C_{D}$ and the average Nusselt number $\mkern 1.5mu\overline{\mkern-1.5muNu\mkern-1.5mu}\mkern 1.5mu$ for $Re$=10, 20 and 40. As can be seen from the figures and table, the length of the recirculation region and the average Nusselt number depend directly on the value of the Reynolds number, while the drag coefficient varies inversely. A quantitative comparison is also carried out in table 8. Computational results are compared with reference data [30, 31, 33, 37, 34, 36, 28, 35, 8, 38, 17, 29, 32]. Once again, the efficiency of the present scheme is evident from the proximity of our results with the well-established results available in the literature.
@@
Next, we are interested in studying the periodic flow for $Re$=100, 140, and 200. The transition of the vorticity and isotherm contours from steady to periodic state for $Re$=100 is depicted in fig. 17. This may be considered as the representation of the flow for the other $Re$ values considered, although each requires a different time to attain periodicity. As can be seen in fig. 17, no vortex structure formed in the flow field, and the heated cylinder generates a thermal boundary layer near its surface at the beginning. As time marches, two opposite symmetrical vortices are formed simultaneously which grow in width and length with time, and remain symmetrically stable for a certain period as shown in figs. 17 and 17. During this period, the thickness of the thermal boundary layer increases; the temperature distribution also remains symmetric about the line $\theta=0$. After a certain point of time, a fluctuation develops in the flow which destroys the symmetry in the vortices and temperature distribution and they start to oscillate downstream which can be noticed in fig. 17. As the flow develops further, the vortices at the rear of the cylinder grow in size and ultimately are shed from the cylinder towards downstream while heat is carried away. It is heartening to see that the present scheme could capture the characteristic feature of the periodic flow, the so-called von Kármán vortex street very efficiently. With the vortex shedding taking place, isolated hot fluid clusters are observed in the flow (see fig. 17). Subsequently, figs. 17 and 17 indicate the synchronized variation of flow and temperature fields, where vortex shedding phenomena play a determinant role in the heat transfer at the downstream.
@@
The time history of the drag coefficient $(C_{D})$, lift coefficient $(C_{L})$ and average Nusselt number $(\mkern 1.5mu\overline{\mkern-1.5muNu\mkern-1.5mu}\mkern 1.5mu)$ displayed in figs. 18, 18 and 18 pertinently implies that the newly developed scheme has accurately captured the periodic state for all the Reynolds numbers under consideration. The periodicity is further justified by the single dominating peak for $C_{D}$ and $C_{L}$ in the spectral density analysis (see fig. 18). The distribution of local Nusselt number $(Nu^{0})$ over the surface of the cylinder for different $Re$ values are represented in fig. 19. In fig. 19 we carry out a qualitative comparison between the surface distribution of $Nu^{0}$ for $Re$=200 computed in the present study with those of [45, 46]; which reveals close agreement of our results with those taken from the literature. We then compare our numerical values of the flow parameters $C_{D}$, $C_{L}$, $St$ and $\mkern 1.5mu\overline{\mkern-1.5muNu\mkern-1.5mu}\mkern 1.5mu$ with existing studies in table 9. The streakline from our simulation depicting the well-known von Kármán vortex street for $Re$=140 is presented in LABEL:{fig:P4_streak140} along with the streakline reported in the experimental work of Taneda [47]. From the comparison we note that vortex creation and dissipation as captured in the experimental work is effectively computed by the current formulation. Striking similarity between our numerical results and the existing numerical and experimental results for all the cases indeed verify and validate the scheme.
@@ 5 Conclusion
In this work, we extend the philosophy of recent work on 2D transient convection and diffusion equation on Cartesian coordinates [18] to polar coordinates to tackle non-rectangular geometries. The transformation-free higher-order compact finite difference scheme proposed here can be applied directly on polar grids, unlike most of the schemes in the literature on polar grids that use transformation between physical and computational domains. We have also combined the virtue of nonuniformity with the present scheme. This levitates the efficiency and robustness of the scheme as it acquires the advantage of accumulating and scattering grid points as necessary. The scheme is theoretically third-order accurate in space and second-order accurate in time. It is worthwhile to mention that for linear problems with analytical solution the scheme exhibits a spatial convergence of order four, which is higher than the theoretical value. The robustness of the scheme is examined by applying it to as many as three benchmark problems of fluid flow and heat flow, viz. driven polar cavity, natural convection in a horizontal concentric annulus, and forced convection around a heated stationary cylinder. In order to resolve the Neumann type boundary conditions, we have also introduced one-sided approximations for first order derivatives. For the problem of forced convection around a heated stationary cylinder, both steady and periodic solutions are being accurately captured by the present scheme. Additionally, important aspects such as the von Kármán vortex phenomenon and the effect of vortex shedding on heat transfer are also studied comprehensively for this fluid body interaction problem. The accuracy of the computed solutions is estimated from the results obtained in all the test problems, which are in excellent agreement with the existing results both qualitatively and quantitatively. We have reported the perceived order of convergence for all the variables in flow problems that are not supplemented with analytical solutions. It is heartening to see that our solutions could attain the theoretical order of convergence in both space and time in all cases.
Acknowledgement
The second author is thankful to Science & Engineering Research Board, India for assistance under Core Research Grant (Project File Number: CRG/2022/000668).